unh Find the uns an vector, the homol vochor and the curvahure of the : a sin, 2:0 0 he pon) with parameter θ,
Answer #1

Equation of circle

x^2+y^2=a^2

Differentiate

2x+2yrac{dy}{dx}=0

dy

cot θ

A vector along the tangent direction would be

hat{T}proptohat{imath}-cot hetahat{jmath}

Make its modulus 1 by multiplying sin

hat{T}=sin hetahat{imath}-cos hetahat{jmath}

Similarly, slope of normal is

= tan θ

and hence

hat{N}=cos hetahat{imath}+sin hetahat{jmath}

Curvature general formula

2

For the given curve

dx2

kappa=rac{rac{a^2}{y^3}}{left[1+left ( rac{x}{y} ight )^2 ight]^{3/2}}=rac{1}{a}

as expected.